Magnetic force on a charged particle, moving in a magnetic field is given by
1
Expert's answer
2020-11-20T09:32:52-0500

Magnetic field "\vec{B}" acts on a particle with a charge "q" moving with a velocity "\vec{v}" with a force given by Lorentz law :

"\vec{F}=q [\vec{v}\times\vec{B}]" , where I note "[\vec{a}\times\vec{b}]" - vector product of "\vec{a},\vec{b}" .

However in general case we need to consider both magnetic and electrical force on a particle, as these two fields are closely tied on to the other. This expression is called 'generalized Lorentz law' :

"\vec{F}=q [\vec{v}\times\vec{B}]+q\vec{E}"